# Disentangled Source-Free Personalization for Facial Expression Recognition with Neutral Target Data

### v1

@@ 4.4 Visualization of Feature Distributions
5 Conclusion
This paper introduces a novel subject-based SFDA framework that specializes in FER in health monitoring applications in the practical case when only neutral expression control target data is available for adaptation. Unlike conventional SFDA settings, our approach effectively adapts a pre-trained source model to target subjects by leveraging a disentangled generative strategy. We propose a one-stage Disentangled SFDA (DSFDA) framework that integrates data generation with domain adaptation, improving model generalization while mitigating the practical constraints of FER data collection in real-world clinical environments. Experimental results on the BioVid and UNBC-McMaster pain assessment datasets show that DSFDA can significantly outperform SOTA SFDA methods and highlight its effectiveness in handling cross-domain variability and subject-specific adaptations.
Acknowledgement
This work was supported by the Fonds de recherche du Québec – Santé (FRQS), the Natural Sciences and Engineering Research Council of Canada (NSERC), the Canada Foundation for Innovation (CFI), and the Digital Research Alliance of Canada.
@@ Acknowledgement
  • Experimental Results of the Two-Stage Framework: Comprehensive evaluation and analysis of the framework’s performance.
  • Impact of prototype selection methods: Source pain prototype selection using different clustering methods.
Figure 6: The visual results of generated pain images using the ImaGINator model on the BioVid dataset.
@@ Appendix A Two-Stage Methodology
The architecture of the proposed disentangled SFDA (DSFDA) method is presented in the main paper. Algorithm 1 summarizes the DSFDA training process.
@@
Algorithm 1 DSFDA Training Process.
1:Input:
2:$\mathbf{I}_{\text{id}}$: target neutral identity
3:$\mathbf{I}_{\text{exp}}$: target neutral expression
4:$C_{\text{exp}}$: pre-trained expression classifier
5:$D_{\text{id}}$: identity discriminators
6:$D_{\text{exp}}$: expression discriminator
7:$E_{\text{id}}$: identity encoder
8:$E_{\text{exp}}$: expression encoder
9:$D$: decoder
10:Hyperparameters: $\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$, $\lambda_{4}$
11:Output: Adapted generator $G$, Adapted expression classifier $C_{\text{exp}}$
12:Initialization: Generator $G$, Discriminators $D_{\text{id}}$, $D_{\text{exp}}$, Classifier $C_{\text{exp}}$
13:for $t=1:T$ do
14:     Train Generator:
15:     for each pair of images $\mathbf{I}_{\text{id}},\mathbf{I}_{\text{exp}}$ do
16:         Extract features $\mathbf{z}_{\text{id}}=E_{\text{id}}(\mathbf{I}_{\text{id}})$, $\mathbf{z}_{\text{exp}}=E_{\text{exp}}(\mathbf{I}_{\text{exp}})$
17:         Generate fake image $\mathbf{I}_{\text{g}}=G(\mathbf{z}_{\text{id}},\mathbf{z}_{\text{exp}})$
18:         Generate reconstructions $\hat{\mathbf{I}}_{\text{id}},\hat{\mathbf{I}}_{\text{exp}}$
19:     end for
20:     Train Discriminators: Update $\mathcal{L}_{\text{dis\_id}}$, $\mathcal{L}_{\text{dis\_exp}}$
21:     Train Classifier: Apply cross-entropy loss $\mathcal{L}_{\text{exp}}$
22:     Update Generator: Minimize generator loss $\mathcal{L}_{\text{gen}}$
23:end for
24:Return: Adapted expression classifier $C_{\text{exp}}$
Figure 7: The visual results of generated pain images using the FExGAN-Meta model on the BioVid dataset.
@@
Appendix B Experimental results
B.1 Qualitative Results
@@
The visual results of the generated images using ImaGINator and FExGAN-Meta are illustrated in Fig.s 6 and  7, respectively. In addition to the visual results, we compare them using numerical results using three metrics, including structural similarity index measure (SSIM), peak signal-to-noise ratio (PSNR), and fréchet inception distance (FID). SSIM is used as a measure to evaluate the quality of generated images using generative models compared to the ground truth or target images. PSNR is used to assess how closely the generated images match the ground truth images. FID measures the similarity between the distributions of feature representations of real and generated images. In Fig. 8, the generated pain images using our Dis-SFDA method are shown.
B.2 Quantitative Results
@@
The quantitative performance of the generative models used to produce synthetic pain images is presented in Table 4. ImaGINator achieves superior results compared to FExGAN-Meta, with higher SSIM (0.82 vs. 0.69) and PSNR (23.00 vs. 22.14), as well as a lower FID score (99.50 vs. 110.58), indicating that ImaGINator generates more realistic and high-quality images, resulting in improved adaptation performance.
@@
We compared the results for the same settings on the UNBC-McMaster dataset using the two-stage framework in Table 5. In the Source Only setting, the pre-trained source model achieves an average accuracy of 61.30%. Using synthetic pain images in the Org. Neutral & Syn. Pain setting improves the results of adaptation significantly compared to the Source Only setting. In this category, SHOT (68.27%) and NRC (68.37%) show comparable results, while the proposed method, SFDA-DE (ours), outperforms both with an average accuracy of 70.41%, demonstrating the effectiveness of our approach. Finally, in the Oracle setting, where the model is fine-tuned on labeled target data, the framework achieves the highest overall accuracy of 97.59%, setting the performance ceiling under ideal conditions.
Figure 8: The visual results of generated pain images using our Dis-SFDA model on the BioVid dataset.
Appendix C Impact of prototype selection methods.
We evaluated the effectiveness of various image prototype selection strategies from the source domain by comparing random, match, DBSCAN, and $k$-means clustering methods and their impact on adaptation performance, as shown in Fig. 9. For DBSCAN, we varied the hyperparameters epsilon (the maximum distance between two points for them to be considered neighbors) and min_samples (the minimum number of points required to form a dense region). These changes result in different clustering outcomes. In DB-All and DB-Sub, the algorithm was applied to all 77 subjects collectively and each subject individually. In DB-D, the most dense cluster for each subject was selected, while in DB-Cls, we selected the cluster closest to the target subject’s cluster. For the $K$-means we used two configurations: $K$=$77$ (KM-All), where the closest point to the centroid was chosen for all subjects, and $K$=$1$ (KM-Sub), where the closest point to the centroid was selected for each subject individually. Based on the results, subject-specific clustering significantly improved performance in both cases. Applying DBSCAN to individual subjects outperformed the all-subject approach by 3.72%, while for $K$-means, the improvement was 4.83%. Based on the average results, utilizing matched videos from the source domain significantly enhances performance more than using a random video. This improvement can be attributed to the shift in expression between different subjects; each individual expresses pain in distinct ways and with varying intensity. For instance, facial cues may be more pronounced in women and younger subjects compared to men and older subjects. Consequently, the model can better disentangle expression from other information when using videos that align with the age and gender of the target subjects.
Table 4: Quantitative results of the generative models on BioVid dataset.
@@
Table 5: Accuracy (%) of the proposed method and state-of-the-art methods on the UNBC-McMaster dataset for 5 target subjects with all 20 sources. Bold text shows the highest accuracy.

### v4

@@ IV-D Visualization of Feature Distributions
V Conclusion
This paper introduces a subject-based SFDA method for FER in real-world health diagnosis and monitoring applications from Q&A videos. This method is specialized for the practical case where only neutral expression control target data is available for adaptation. Unlike conventional SFDA settings, our approach effectively adapts a pre-trained source model to target subjects by leveraging a disentangled generative strategy. We propose a one-stage disentangled SFDA (DSFDA) framework that integrates data generation with domain adaptation, improving model generalization while mitigating the practical constraints of FER data collection in real-world clinical environments. Experimental results on the BioVid and UNBC-McMaster pain assessment and on the StressID datasets show that DSFDA can significantly outperform SOTA SFDA methods. Results highlight its effectiveness in handling cross-domain variability and subject-specific adaptations.
Acknowledgments: This work was supported in part by the Fonds de recherche du Quebec – Santé (FRQS), the Natural Sciences and Engineering Research Council of Canada (NSERC), Canada Foundation for Innovation (CFI), and the Digital Research Alliance of Canada.
VI Ethical Impact Statement
Potential harms to human subjects. This study did not involve new data collection. It utilizes the publicly available BioVid Heat Pain dataset [8] to evaluate the proposed method. Access to the BioVid dataset requires a request to the authors of the dataset. The dataset includes participants from three age groups: (1) 18–35, (2) 36–50, and (3) 51–65, with an equal distribution of male and female subjects in each group, minimizing potential age and gender bias. All participants, except for 13 subjects, provided informed consent for their data (including images and videos) to be used for research dissemination through online publications, conference presentations, and proceedings. In this paper, the data from these 13 participants were not used to illustrate research results.
Additionally, we used the UNBC-McMaster Shoulder Pain dataset [59], which requires submitting a request to the authors and signing an access agreement. We also employed the StressID dataset [62], which includes video, audio, and physiological recordings collected under ethical approval. Access to the StressID dataset requires signing a data usage agreement, and it is limited to non-commercial scientific research. All participants provided informed consent, and data usage must comply with GDPR and privacy guidelines. Only the visual modality (video recordings) was used in this study.
Potential negative societal impact. The generation of pain images has the potential to be misused in order to create false evidence of pain or medical conditions in healthcare or legal settings. Pain images could be exploited to influence emotions in contexts such as fundraising campaigns or social media, where the aim may be to elicit sympathy or alarm from the audience in order to further a particular agenda.
Limits of generalizability. The proposed method has been validated for neutral and pain levels using the BioVid, UNBC-McMaster, and StressID datasets. However, we do not claim that the method is generalizable to all emotions or stress conditions, as further validation is necessary to confirm its effectiveness across a broader range of expressions and emotional states.
Risk-mitigation strategies. To mitigate the misuse risks, licensing and secure authentication protocols to restrict access to the generative model is an effective strategy. This approach ensures that only authorized users can access the models, while strict usage policies help prevent misuse. Moreover, authorities can detect and identify fake images using common and widely available tools, providing an additional layer of protection against potential misuse. While we will make our code publicly available, the pre-trained weights will not be released and will only be provided upon request from trusted parties. We will provide an ethical impact statement with the paper to mitigate the risk of misunderstanding about our method. Additionally, we will explicitly clarify which emotions this work has been validated for, emphasizing that its effectiveness beyond these validated emotions requires further investigation and validation.
@@ VI Ethical Impact Statement
  • Ablation Studies: Ablation studies on image generation methods.
@@ Appendix A Two-Stage Methodology
Figure 6: Visual comparison of generated pain images produced by different models: (a) ImaGINator, (b) FExGAN, and (c) DSFDA.
Appendix B Ablation Studies
@@
Ablation on the generation results. The visual results of the generated images using ImaGINator and FExGAN-Meta are illustrated in Fig.s 6(a) and  6 (b), respectively. In addition to the visual results, we compare them using numerical results using three metrics, including structural similarity index measure (SSIM), peak signal-to-noise ratio (PSNR), and fréchet inception distance (FID). SSIM is used as a measure to evaluate the quality of generated images using generative models compared to the ground truth or target images. PSNR is used to assess how closely the generated images match the ground truth images. FID measures the similarity between the distributions of feature representations of real and generated images. In Fig. 6 (c), the generated pain images using our DSFDA method are shown.
@@
The quantitative performance of the generative models used to produce synthetic pain images is presented in Table VI. ImaGINator achieves superior results compared to FExGAN-Meta, with higher SSIM (0.82 vs. 0.69) and PSNR (23.00 vs. 22.14), as well as a lower FID score (99.50 vs. 110.58), indicating that ImaGINator generates more realistic and high-quality images, resulting in improved adaptation performance.
@@
We compared the results for the same settings on the UNBC-McMaster dataset using the two-stage framework in Table VII. In the Source Only setting, the pre-trained source model achieves an average accuracy of 61.30%. Using synthetic pain images in the Org. Neutral & Syn. Pain setting improves the results of adaptation significantly compared to the Source Only setting. In this category, SHOT (68.27%) and NRC (68.37%) show comparable results, while the proposed method, SFDA-DE (ours), outperforms both with an average accuracy of 70.41%, demonstrating the effectiveness of our approach. Finally, in the Oracle setting, where the model is fine-tuned on labeled target data, the framework achieves the highest overall accuracy of 97.59%, setting the performance ceiling under ideal conditions.
TABLE VI: Quantitative results of the generative models on BioVid dataset.
@@
Impact of prototype selection methods. We evaluated the effectiveness of various image prototype selection strategies from the source domain by comparing random, match, DBSCAN, and $k$-means clustering methods and their impact on adaptation performance, as shown in Fig. 7. For DBSCAN, we varied the hyperparameters epsilon (the maximum distance between two points for them to be considered neighbors) and min_samples (the minimum number of points required to form a dense region). These changes result in different clustering outcomes. In DB-All and DB-Sub, the algorithm was applied to all 77 subjects collectively and each subject individually. In DB-D, the most dense cluster for each subject was selected, while in DB-Cls, we selected the cluster closest to the target subject’s cluster. For the $K$-means we used two configurations: $K$=$77$ (KM-All), where the closest point to the centroid was chosen for all subjects, and $K$=$1$
(KM-Sub), where the closest point to the centroid was selected for each subject individually. Based on the results, subject-specific clustering significantly improved performance in both cases. Applying DBSCAN to individual subjects outperformed the all-subject approach by 3.72%, while for $K$-means, the improvement was 4.83%. Based on the average results, utilizing matched videos from the source domain significantly enhances performance more than using a random video. This improvement can be attributed to the shift in expression between different subjects; each individual expresses pain in distinct ways and with varying intensity. For instance, facial cues may be more pronounced in women and younger subjects compared to men and older subjects. Consequently, the model can better disentangle expression from other information when using videos that align with the age and gender of the target subjects.
TABLE VII: Accuracy (%) of the proposed method and state-of-the-art methods on the UNBC-McMaster dataset for 5 target subjects with all 20 sources. Bold text shows the highest accuracy.
@@
Figure 8: Comparison of pseudo-label generation reliability between SHOT and NRC methods using FExGAN and ImaGINator to generate missing classes.
Impact of generation methods on pseudo-labels. We evaluated the reliability of pseudo-labels produced by SHOT and NRC methods when using FExGAN and ImaGINator to generate the missing classes in the target data. ImaGINator consistently produced a higher ratio of reliable pseudo-labels compared to FExGAN. Specifically, the NRC method achieved 75% reliability with FExGAN and improved to 78% with ImaGINator, while the SHOT method resulted in 70% with FExGAN and reached 80% with ImaGINator. Notably, the combination of SHOT with FExGAN yielded the lowest pseudo-label ratio (70%), suggesting that SHOT is less effective when paired with FExGAN compared to ImaGINator. Furthermore, the NRC-FExGAN configuration outperformed our DSFDA-FExGAN method, likely because the generated pain images – often resembling neutral expressions – received reliable pseudo-labels. Since most real or generated pain images do not visibly express pain, assigning pain labels indiscriminately can introduce misleading representations, ultimately diminishing the model’s performance during testing.